OpenStudy (pokiedokie):

3(x+1) = 9+2x i know i'm probably making this harder than it actually is but i'm doubting if how i'm solving it is right?? Am i supposed to distribute 3 to (x+1)?? (sorry, i'm bad at math)

OpenStudy (anonymous):

3x+3=9+2x yes you are then simplify and get x alone

OpenStudy (pokiedokie):

thank you!! and for 9+2x all i do is divide correct?

OpenStudy (lynfran):

no

OpenStudy (lynfran):

3x+3=9+2x subtract 2x frm both sides we get 3x-2x+3=9+2x-2x 3x-2x+3=9 now subtract 3 frm both sides and simplify

OpenStudy (pokiedokie):

thank you very much, i was confused but now i think i get it a little better

OpenStudy (lynfran):

welcome